OpenStudy (anonymous):

Soybean is 16% protien, cornmeal is 8% protien.How many pounds of each should be mixed together in order to get a 320 pound mixture that is 12% protien?

OpenStudy (saifoo.khan):

.16s+.08(320-s)=.12*320 s = 160 s=160 pounds of soybeans are used. 320-160=160 pounds of cornmeal are used.

OpenStudy (saifoo.khan):

let me prove.

OpenStudy (saifoo.khan):

Proof: .16*160+.08*160=.12*320 12.8+19.2=32 32=32

OpenStudy (anonymous):

anyone disagree with saifoo?

OpenStudy (saifoo.khan):

you? lol

OpenStudy (anonymous):

alrighty then, thanks for your efforts.

OpenStudy (anonymous):

we're about to be shut down any way.

OpenStudy (saifoo.khan):

but we will be back.

OpenStudy (saifoo.khan):

2mins.